Calculate the value of induced emf between the ends of an axle of a railway carriage
1.75 m long traveling on level ground with a uniform velocity of 50 km per hour. The
vertical component of Earth’s magnetic field (Br) is given to be 5  105T.
1
Expert's answer
2020-08-24T13:24:33-0400

"Given\\L=1.75m\\v=50\frac{km}{h}=\frac{50000m}{3600c}=13.89\frac{m}{c}\\B_r=5\times10^{-5}T;\\Solution\\E=B_rLv(sin\alpha)=1.215\times10^{-3}V(sin90=1)\\Answer:E(emf)=1.215mV"